PLEASE ANSWER IN GENERAL TERMS. EXPLAIN IT TO ME LIKE I AM 5 YEARS OLD.

1. What is the apparent rule for combining equal resistances in series circuit? In parallel circuits?

2. What is the apparent rule for combining unequal resistances in series circuits? In parallel circuits?

3. What is the apparent rule for the total resistance when resistors are added up in series? In parallel?
Answer #1

1. The rule is that if the if n equal resistances R each,  are in series, they can be combined into one with resistance = n x R

For parallel = R/n

2.

The rule is that if the if n unequal resistances R1 , R2, R3,... Rn,  are in series, they can be combined into one with resistance = R1+R2+R3....Rn

For parallel =(1/R1+1/R2+1/R3....1/Rn)-1

3.

The rule is that if the if n unequal resistances R1 , R2, R3,... Rn,  are in series, they can be combined into one with resistance = R1+R2+R3....Rn

For parallel =(1/R1+1/R2+1/R3....1/Rn)-1